Casper Inc. buys and sells a household cleaning item for the last 8 years at their New
Jersey outlet.
The following information is for Casper Inc. for the month ended June 30 2018. Casper
Inc. uses the periodic method for inventory.
Date Description Quantity Unit Cost or Selling

Price
June 1 Opening Inventory 80 80
June 3 Purchase 270 86
June 5 Sale 220 140
June 9 Sale 15 25
June 11 Sales Return 30 140
June 13 Purchase 110 92
June 15 Purchase Return 20 92
June 17 Sale 130 152
June 19 Purchase 70 100

Required:
a. Calculate
(i) Ending Inventory,
(ii) Cost of Goods sold,
(iii) Gross Profit
Under each of the following methods. LIFO. (2) FIFO. (3) Average-cost.

b. From your calculations, analyze which method results in
1. The highest ending inventory, and
2. The highest cost of goods sold.
Answer #1

LIFO

Date Particulars Receipts Issues Balance
Qty Rate Amount Qty Rate Amount Qty Rate Amount
June1 Opening Invetory 80 80 6400 0 0 0 80 80 6400
June3 Purchase 270 86 23220 0 0 0 80 80 6400
270 86 23220
June5 Sales 0 0 0 220 140 30800 80 80 6400
50 86 4300
June9 Sales 0 0 0 15 25 375 80 80 6400
35 86 3010
June11 Sales Returns 30 140 4200 0 0 0 80 80 6400
65 86 5590
June13 Purchase 110 92 10120 0 0 0 80 80 6400
65 86 5590
110 92 10120
June15 Purchase Return 0 0 0 20 92 1840 80 80 6400
65 86 5590
90 92 8280
June17 Sales 0 0 0 130 152 19760 80 80 6400
25 86 2150
June19 Purchase 70 100 7000 0 0 0 80 80 6400
25 86 2150
70 100 7000
June3 Closing Inventry 175 15550

LIFO

1. Inventry = 15550

2. Cost of Goods Sold=(220+15)*86=20210 +(90*92)+(40*86)=31930

3. Gross Profit=Sales-COGS

GP=50935-31930

=19005